OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

\[Given~that~3^{2x}-1-7=0,\]find the value of x by giving the answer correct to 2 decimal places.

OpenStudy (anonymous):

\[3^{2x}=8\]

OpenStudy (anonymous):

take the natural logarithm of both sides

OpenStudy (anonymous):

\[\log _{10}3^{2x}=\log _{10}8\]

OpenStudy (anonymous):

Yes, that works to, 2x comes down

OpenStudy (anonymous):

\[2xlog _{10}3=\log _{10}8\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[2x=\frac{ \log _{10}8 }{ \log _{10}3 }\]

OpenStudy (anonymous):

Yup, finish it off

OpenStudy (anonymous):

\[x=0.95\]

OpenStudy (anonymous):

Thnx @iambatman :)

OpenStudy (anonymous):

Np